OpenStudy (logan13):

Variable y varies directly with variable x, and y = 45 when x = 9. What is value of y when x = 2? y =

OpenStudy (qwertty123):

Direct variation means their ratio is a constant x/y is 9/45=1/5 so when x=2, y=5*2=10

OpenStudy (logan13):

10

OpenStudy (qwertty123):

Yup

OpenStudy (logan13):

thanks

OpenStudy (qwertty123):

You welcome! Sorry It took me a while to reply!

OpenStudy (logan13):

its fine